# CAD: Confidence-Aware Adaptive Displacement for Semi-Supervised Medical Image Segmentation

### v1

@@ Abstract
Semi-supervised medical image segmentation aims to leverage minimal expert annotations, yet remains confronted by challenges in maintaining high-quality consistency learning. Excessive perturbations can degrade alignment and hinder precise decision boundaries, especially in regions with uncertain predictions. In this paper, we introduce Confidence-Aware Adaptive Displacement (CAD), a framework that selectively identifies and replaces the largest low-confidence regions with high-confidence patches. By dynamically adjusting both the maximum allowable replacement size and the confidence threshold throughout training, CAD progressively refines the segmentation quality without overwhelming the learning process. Experimental results on public medical datasets demonstrate that CAD effectively enhances segmentation quality, establishing new state-of-the-art accuracy in this field. The source code will be released after the paper is published.
@@ I Introduction
In medical image segmentation, semi-supervised medical image segmentation (SSMIS) methods have emerged to enhance model performance by utilizing a small amount of labeled data and a large amount of unlabeled data[2, 3, 4] . This approach alleviates the challenge of limited labeled data and effectively leverages the rich information contained in the unlabeled dataset. In recent years, as this field has advanced, the performance of SSMIS has increasingly approached that of fully supervised segmentation methods, particularly under conditions with minimal labeled data, showing remarkable segmentation results[5, 6]. This progress has been driven by several techniques, with consistency learning being one of the core methods in semi-supervised learning[7, 8, 9]. Consistency learning facilitates the effective use of unlabeled data by enforcing consistent predictions across different augmentations of the input data, utilizing pseudo-labels and data augmentation strategies to generate high-quality labels[10]. Common techniques in consistency learning include self-training and consistency loss, which minimize the discrepancy between model outputs under different augmentations, making the model more stable and reliable when processing unlabeled data[11, 12].
In consistency learning, perturbations such as data augmentations and feature map perturbations play a pivotal role[12]. These perturbations ensure that the model learns robust features by making the model’s predictions consistent despite variations in the input data. Data augmentations, including rotation, scaling, and flipping, challenge the model, forcing it to focus on stable and generalizable features while ignoring spurious ones[13, 14]. On the other hand, feature map perturbations, such as adding noise or applying dropout, directly alter the feature representations, encouraging the model to maintain consistency even in the presence of noisy or incomplete information[10, 15, 16]. Together, these perturbations help the model refine its internal representations, ultimately enhancing its performance on unlabeled data and improving segmentation accuracy in semi-supervised settings.
@@ II Related Works
Consistency learning plays a pivotal role in SSMIS by utilizing the large amounts of unlabeled data available. The core idea is to encourage the model to produce stable predictions under various perturbations or augmentations of the input data, improving performance even with limited labeled data. Early works, such as the study by Sajjadi et al. [7], highlighted the importance of stochastic transformations, like dropout and random augmentations, to regularize deep networks and improve generalization, making the model more robust to input variability. More recent studies have built on these techniques to enhance consistency in semi-supervised segmentation tasks. For instance, ConMatch[18] incorporates confidence-guided consistency regularization, which improves performance by refining pseudo-label confidence estimation.
A key challenge in consistency learning is ensuring the model learns meaningful features without being overly sensitive to noisy predictions. To address this, several works have introduced auxiliary methods or refined model predictions. For example, the UDiCT[19], pairs annotated and unannotated data based on uncertainty to mitigate the impact of unreliable pseudo-labels. Similarly, Huang et al. [17] introduced a two-stage approach for segmenting neurons from electron microscopy volumes, improving model robustness by enforcing consistency across perturbed versions of unlabeled data. Another notable contribution is from Tarvainen and Valpola [9], who introduced the Mean Teacher model, which maintains an exponential moving average (EMA) of model weights to improve segmentation performance in semi-supervised settings.
In medical image segmentation, where data variability presents significant challenges, consistency learning has proven effective in overcoming the limitations of insufficient labeled data. Methods like the one proposed by Shu et al. [20] and Wang et al. [21] have explored augmenting consistency learning with additional tasks or uncertainty estimation. For example, Liu et al. [22] applied transformer-based models for COVID-19 lesion segmentation, using consistency across augmented views to address the shortage of labeled data. Furthermore, Chen et al. [23] proposed Cross Pseudo Supervision (CPS), where two segmentation networks with different initializations enforce consistency on each other’s predictions, thereby expanding the training data with pseudo-labels. These advances demonstrate how consistency learning, particularly when combined with novel architectures, can significantly improve model performance in medical image segmentation with limited labeled data.
@@ III-B1 Mean Teacher Loss ($L_{\text{MT}}$)
For labeled data, the Mean Teacher loss is given by:
$L_{\text{MT}}=L_{\text{dice}}+L_{\text{ce}},$ (11)
where $L_{\text{dice}}$ is the Dice loss and $L_{\text{ce}}$ is the Cross Entropy loss. The Dice loss is defined as:
$L_{\text{dice}}=1-\frac{2\sum_{c}\text{true}_{c}\cdot\hat{\text{pred}}_{c}}{\sum_{c}\text{true}_{c}+\sum_{c}\hat{\text{pred}}_{c}},$ (12)
where $\text{true}_{c}$ and $\hat{\text{pred}}_{c}$ are the ground truth and predicted probabilities for class $c$, respectively.
The Cross Entropy loss is computed as:
$L_{\text{ce}}=-\sum_{c}\text{true}_{c}\log(\hat{\text{pred}}_{c}),$ (13)
where $\hat{\text{pred}}_{c}=\text{softmax}(\text{logits}_{c})$ is the predicted probability for each class.
@@ III-B3 Confidence-Aware Displacement Loss ($L_{\text{CAD}}$)
By incorporating the CAD loss, the model is encouraged to refine its predictions and improve the segmentation by focusing on uncertain regions and using high-confidence predictions from the other model to guide the correction of low-confidence areas.
@@ IV-A Datasets
In this study, we evaluate the proposed semi-supervised learning approach using two widely recognized 3D medical image datasets: the ACDC (Automated Cardiac Diagnosis Challenge)[24] dataset and the PROMISE12[25] dataset. While both datasets originally consist of 3D images, we perform our experiments using their corresponding 2D slices for training and evaluation. The 2D slices provide a simplified approach to segmentation, which aligns with common practices in the field[26].
@@ IV-A1 ACDC Dataset
The ACDC dataset[24] consists of 200 annotated short-axis cardiac cine-MR images from a cohort of 100 patients, each containing four distinct classes: left ventricle, right ventricle, myocardium, and background. For the purpose of this study, the 3D scans were converted into 2D slices, with each slice representing a cross-sectional view of the heart. The dataset is split into 70 patients’ scans for training, 10 for validation, and 20 for testing.
@@ IV-A2 PROMISE12 Dataset
The PROMISE12 dataset[25], created for the MICCAI 2012 prostate segmentation challenge, contains MRI scans from 50 patients with various prostate diseases. Each of the original 3D scans is converted into a series of 2D slices, which we use for the segmentation task. These slices present challenges due to the variations in prostate shape, size, and other characteristics across different patients.

### v2

@@ Abstract
Semi-supervised medical image segmentation aims to leverage minimal expert annotations, yet remains confronted by challenges in maintaining high-quality consistency learning. Excessive perturbations can degrade alignment and hinder precise decision boundaries, especially in regions with uncertain predictions. In this paper, we introduce Confidence-Aware Adaptive Displacement (CAD), a framework that selectively identifies and replaces the largest low-confidence regions with high-confidence patches. By dynamically adjusting both the maximum allowable replacement size and the confidence threshold throughout training, CAD progressively refines the segmentation quality without overwhelming the learning process. Experimental results on public medical datasets demonstrate that CAD effectively enhances segmentation quality, establishing new state-of-the-art accuracy in this field.
@@ I Introduction
In medical image segmentation, semi-supervised medical image segmentation (SSMIS) methods have emerged to enhance model performance by utilizing a small amount of labeled data and a large amount of unlabeled data[2, 3]. This approach alleviates the challenge of limited labeled data and effectively leverages the rich information contained in the unlabeled dataset. In recent years, as this field has advanced, the performance of SSMIS has increasingly approached that of fully supervised segmentation methods, particularly under conditions with minimal labeled data, showing remarkable segmentation results[4, 5]. This progress has been driven by several techniques, with consistency learning being one of the core methods in semi-supervised learning[6, 7, 8], which facilitates the effective use of unlabeled data by enforcing consistent predictions across different augmentations of the input data, utilizing pseudo-labels and data augmentation strategies to generate high-quality labels[9]. Common techniques in consistency learning include self-training and consistency loss, which minimize the discrepancy between model outputs under different augmentations, making the model more stable and reliable when processing unlabeled data[10, 11]. Within consistency learning, perturbations such as data augmentations and feature map perturbations play a pivotal role[11], ensuring that the model learns robust features by making predictions consistent despite variations in the input data. Data augmentations, including rotation, scaling, and flipping, challenge the model, forcing it to focus on stable and generalizable features while ignoring spurious ones[12, 13]; meanwhile, feature map perturbations, such as adding noise or applying dropout, directly alter the feature representations, encouraging the model to maintain consistency even in the presence of noisy or incomplete information[9, 14]. Together, these perturbations help the model refine its internal representations, ultimately enhancing its performance on unlabeled data and improving segmentation accuracy in semi-supervised settings.
@@ II Related Works
Consistency learning plays a pivotal role in SSMIS by effectively utilizing large amounts of unlabeled data. The core idea is to encourage models to produce stable predictions under various perturbations or augmentations of the input data, thereby enhancing performance even with limited labeled examples. Early works, such as the study by Sajjadi et al. [6], highlighted the importance of stochastic transformations like dropout and random augmentations to regularize deep networks, improving generalization and robustness against input variability. More recent studies have expanded on these ideas to enhance consistency in semi-supervised segmentation tasks; for instance, ConMatch[16] incorporates confidence-guided consistency regularization, refining pseudo-label confidence estimations to improve performance. Additionally, methods such as UDiCT[17] pair annotated and unannotated data based on uncertainty, thereby mitigating the impact of unreliable pseudo-labels, while Huang et al. [15] introduced a two-stage approach enforcing consistency across perturbed versions of unlabeled electron microscopy volumes to enhance model robustness.
In medical image segmentation, consistency learning addresses significant challenges posed by data variability and insufficient labeled data. Prominent methods have augmented consistency learning by integrating additional tasks or uncertainty estimations, as demonstrated by Shu et al. [18] and Wang et al. [19]. For instance, Liu et al. [20] applied transformer-based models to COVID-19 lesion segmentation, using consistency across augmented views to alleviate the shortage of labeled data. Moreover, Chen et al. [21] proposed Cross Pseudo Supervision (CPS), where two segmentation networks with different initializations enforce consistency on each other’s predictions, effectively expanding training data via pseudo-labels. Another notable contribution is from Tarvainen and Valpola [8], who introduced the Mean Teacher model, maintaining an exponential moving average (EMA) of model weights to improve segmentation performance under semi-supervised conditions. Collectively, these advances demonstrate how consistency learning, especially when combined with innovative model architectures and training strategies, significantly enhances segmentation performance with limited labeled data.
@@ III-B1 Mean Teacher Loss ($L_{\text{MT}}$)
For labeled data, the Mean Teacher loss combines the Dice loss and the Cross Entropy loss, given by:
$L_{\text{MT}}=\left[1-\frac{2\sum_{c}\text{true}_{c}\cdot\hat{\text{pred}}_{c}}{\sum_{c}\text{true}_{c}+\sum_{c}\hat{\text{pred}}_{c}}\right]-\sum_{c}\text{true}_{c}\log(\hat{\text{pred}}_{c}),$ (11)
where $\text{true}_{c}$ and $\hat{\text{pred}}_{c}=\text{softmax}(\text{logits}_{c})$ represent the ground truth and the predicted probability for class $c$, respectively.
@@ IV-A Datasets
In this study, we evaluate the proposed semi-supervised learning approach using two widely recognized datasets: ACDC (Automated Cardiac Diagnosis Challenge)[22] and PROMISE12[23]. Although originally composed of 3D medical images, both datasets are converted into 2D slices for simplicity, aligning with common practices in segmentation studies[24].
@@ IV-A1 ACDC Dataset
The ACDC dataset[22] contains 200 short-axis cardiac cine-MR images from 100 patients, labeled into four classes: left ventricle, right ventricle, myocardium, and background. Each 3D scan is converted into 2D cross-sectional slices, and the dataset is partitioned into training, validation, and testing sets with a ratio of 70:10:20.
@@ IV-A2 PROMISE12 Dataset
The PROMISE12 dataset[23], designed for the MICCAI 2012 prostate segmentation challenge, includes MRI scans from 50 patients with diverse prostate pathologies. We similarly convert the original 3D images into 2D slices, which pose segmentation challenges due to substantial variations in prostate size and shape among patients.